OpenStudy (anonymous):

Simplify, leave in factored form: 5y/y(4y-1)+ 2/4y-1

OpenStudy (jusaquikie):

\[\frac{ 5y }{y(4y-1)}+\frac{ 2 }{ 4y-1}\]

OpenStudy (anonymous):

Yes,

OpenStudy (jusaquikie):



OpenStudy (jusaquikie):



OpenStudy (jusaquikie):

not sure where the 4 came from lol it's should be 5+2 = 7 / 4y-1 it's late lol

OpenStudy (anonymous):

Phew, I thought I was losing my mind!

OpenStudy (anonymous):

Thank you very much!

OpenStudy (jusaquikie):

np good night